.03x+.05(3x)=5400
this is the equation i came up with from a word problem trying to figure out how much was invested at 3% annual interest and, 3 times that at 5% interest if the yearly total of both is 5400.
I don't know what to do next though

Answer by rfer(16322)

.05*3x+.03x=5400
.15x+.03x-162=5400
.18x+5400
x=30000
.03x=900
.05x*3=4500